# First Measurement of Correlated Charge Noise in Superconducting Qubits at an Underground Facility 32pt32pt

### v1

@@
We measure space- and time-correlated charge jumps on a four-qubit device, operating 107 meters below the Earth’s surface in a low-radiation, cryogenic facility designed for the characterization of low-threshold particle detectors. The rock overburden of this facility reduces the cosmic ray muon flux by over compared to laboratories at sea level. Combined with coverage of a movable lead shield, this facility enables quantifiable control over the flux of ionizing radiation on the qubit device. Long-time-series charge tomography measurements on these weakly charge-sensitive qubits capture discontinuous jumps in the induced charge on the qubit islands, corresponding to the interaction of ionizing radiation with the qubit substrate. The rate of these charge jumps scales with the flux of ionizing radiation on the qubit package, as characterized by a series of independent measurements on another energy-resolving detector operating simultaneously in the same cryostat with the qubits. Using lead shielding, we achieve a minimum charge jump rate of mHz, almost an order of magnitude lower than that measured in surface tests, but a factor of roughly eight higher than expected based on reduction of ambient gammas alone. We operate four qubits for over 22 consecutive hours with zero correlated charge jumps at length scales above three millimeters.
@@
A growing body of evidence indicates that ionizing radiation impacts the coherence of superconducting qubits. Correlation has been measured between the flux of ionizing radiation and the qubit energy relaxation rate () [1, 2]. Ionizing events in a chip substrate cause simultaneous errors in multi-qubit processors [3]. Moreover, the presence of correlated qubit errors, and the rate at which they occur in unshielded laboratories, can interfere with the efficacy of error-correcting surface codes [4, 5]. Researchers have observed charge and parity errors, correlated in time and space, caused by environmental gammas and cosmic rays interacting with the qubit substrate [6, 7, 8].
@@
Here we describe a continuation of the work performed in Ref. [7], using the same four-qubit chip. The charge environment in the qubit substrate responds to ionizing events through multiple physics channels (including e-h pair production, phonon diffusion, and charge trapping) and on time scales ranging from nanoseconds to hours or even days [9, 10, 3]. Accordingly, the qubits in this work are mildly charge-sensitive transmons () that operate as electrometers, each with a sensing area for electric fields in the substrate of hundreds of square microns. The qubit chip was relocated from the Earth’s surface at Madison, WI to the Northwestern EXperimental Underground Site (NEXUS) at Fermilab in Batavia, IL. Previously, correlated jumps in offset charge, associated with gamma ray and cosmic ray impacts, were observed in this qubit array. Underground, over 99 of cosmic ray muons are shielded by the overburden, creating an environment in which the qubit response to gamma radiation can be studied in isolation. We vary the flux of gamma rays incident on the qubit chip and measure the rate and magnitude of ensuing charge burst events via Ramsey tomography. Other energy-resolving detectors operating simultaneously in NEXUS are used to calibrate the flux and spectrum of ionizing radiation. (This study is in some sense complimentary to that of Ref. [8], in which scintillator detectors provide coincidence information between cosmic ray muon events and qubit errors.)
@@
| Livetime | Q1 | Q2 | Q3 | Q4 | Average | |
| Shield Open | 23.9 hr | |||||
| Shield Closed | 22.1 hr | |||||
| Shield Open (Gammas) | - |
| Q1-Q2 | Q3-Q4 | Q1-Q3 | Q1-Q4 | Q2-Q3 | Q2-Q4 | |
| Separation (m) | 640 | 340 | 3195 | 3330 | 3180 | 3240 |
| Shield Open | ||||||
| Shield Closed |
@@
The average charge jump rates that we measure for the SO and SC data are 0.51 and 0.19 mHz respectively, as shown in Table 1. Therefore, closing the shield reduces the rate of qubit charge bursts by only a factor of 2.7, over eight times less than expected based on the factor of 23 reduction in rate measured in the LMO detector. A possible explanation for this discrepancy is that in our lowest-background SC configuration, we are sensitive to a source of charge bursts that is not dominated by the external gamma flux. The expected muon flux through the qubit (for both configurations) is 0.08 mHz/cm2 – too low to explain the observed rates. (The NuMI muon neutrino beam was not active during data collection for this work. Neutrino interactions in matter [10] therefore contributed no additional muon flux during this study.)
We infer that the ambient gamma flux does not contribute to our SC dataset. We therefore subtract the SC rates from the SO data to obtain a reduced burst rate in the SO data, induced by ambient gammas, of mHz. Based on the LMO data (see Supplemental Material), we estimate our ambient gamma flux in the SO configuration to be approximately five times lower than that measured using a NaI detector in Ref. [7]. If we assume, as Ref. [7] does, that the surface burst rate of 1.35 mHz was gamma-dominated, our ambient rate estimate of mHz is proportional with that expected reduction in flux.
The distances between each qubit pair are different. See Table 2. In the SO data, we measure a correlated charge jump rate in nearby qubit pairs of 0.27 mHz for qubits 1 and 2, and 0.29 mHz for qubits 3 and 4. Given the low single-qubit jump rates, the probability of correlated jumps being due to stochastic coincidence from separate burst events in or near two qubits is orders of magnitude lower than the measured correlated jump rate. This correlated rate is approximately half of the single-qubit rate (to within statistical uncertainty), which is consistent with that observed in Ref. [7] when dominated by gamma flux. In the SO data, the correlated jump rate across distant qubit pairs is too low to make any statements on burst origin with any statistical significance. The same statistical limitation is true for nearby qubit pairs in the SC data. We are able to, for the first time, eliminate correlated charge noise in charge-sensitive qubits separated by over 3 mm, on timescales nearing one day.
@@
The NEXUS DR, shown in Figure S1, is enclosed in a three-part lead shield: (1) a 5.2-inch-thick inner lead plug above the payload, thermalized at 1 K; (2) a 4-inch-thick stationary wall; and (3) a cart-mounted, movable lead shield with a nominal thickness of 4 inches. Taken together, these three shield components provide full, coverage of experimental payloads against ambient gamma radiation. The qubit chip is installed inside of a 1-mm thick A4K hermetic can from Amuneal, thermalized to the DR mixing chamber plate. The DR itself is further equipped with a 1-mm thick A4K magnetic shield from Amuneal, thermalized at 4 K to reduce the magnetic field in the payload region. Finally, a external Metglass blanket surrounds the 300-K shell. The DR base temperature was stabilized at 10.5 mK during the collecation of all data presented here.
@@
Two different warm RF systems are used for qubit characterization and measurement. Both are connected to the qubit input and output ports via a warm RF switch. A vector network analyzer (not shown in Fig. S2) is used for continuous wave measurements during initial qubit characterization. Pulsed RF measurements are performed using the Quantum Instrumentation and Control Kit (QICK): a Xilinx RFSoC board with custom open-source software and firmware for the control of qubit systems [20, 21]. The qubit charge bias voltage is supplied by an arbitrary waveform generator (AWG). The system has since been upgraded to allow for multiplexed simultaneous readout, but at the time of the measurement presented in the main text, this was not possible. As such, individual qubits are measured consecutively for each bias voltage in a Ramsey tomography scan before moving onto the next voltage value.
@@
The efficiency of this method is assessed by applying it to synthetic data with charge jumps inserted at known locations. For each qubit we simulate 1600 scans with a jump rate of 1.1 mHz (slightly higher than the actual rate to ensure we do not see effects of pile-up). The jumps are randomly injected with sizes selected from a flat probability distribution between 0.01 and 0.5. The simulated scans are produced using the templates discussed above, convolved with a Gaussian noise spectrum to add the characteristic noise of our data. The accuracy of the jump-finding algorithm is gauged by the fraction of known charge jumps in the synthetic data that are correctly identified. Systematic errors on this quantity are derived from the standard deviation of this value for 15 different jump sizes across 75 sets of simulated scans per qubit (see Table S2). This effectiveely captures the variation in efficiency across the range of , which can be crudely approximated by Gaussian (with slightly lower efficiency near the threshold in particular).

### v3

@@
We measure space- and time-correlated charge jumps on a four-qubit device, operating 107 meters below the Earth’s surface in a low-radiation, cryogenic facility designed for the characterization of low-threshold particle detectors. The rock overburden of this facility reduces the cosmic ray muon flux by over compared to laboratories at sea level. Combined with coverage of a movable lead shield, this facility enables quantifiable control over the flux of ionizing radiation on the qubit device. Long-time-series charge tomography measurements on these weakly charge-sensitive qubits capture discontinuous jumps in the induced charge on the qubit islands, corresponding to the interaction of ionizing radiation with the qubit substrate. The rate of these charge jumps scales with the flux of ionizing radiation on the qubit package, as characterized by a series of independent measurements on another energy-resolving detector operating simultaneously in the same cryostat with the qubits. Using lead shielding, we achieve a minimum charge jump rate of mHz, almost an order of magnitude lower than that measured in surface tests, but a factor of roughly seven higher than expected based on reduction of ambient gammas alone. We operate four qubits for over 22 consecutive hours with zero correlated charge jumps at length scales above three millimeters.
@@
A growing body of evidence indicates that ionizing radiation impacts the coherence of superconducting qubits. Correlation has been measured between the flux of ionizing radiation and the qubit energy relaxation rate () [1, 2]. Ionizing events in a chip substrate cause simultaneous errors in multi-qubit processors [3]. Moreover, the presence of correlated qubit errors, and the rate at which they occur in unshielded laboratories, can interfere with the efficacy of error-correcting surface codes [4, 5]. Researchers have observed charge and parity errors, correlated in time and space, caused by environmental gammas and cosmic rays interacting with the qubit substrate [6, 7, 8]. Ionizing radiation has also been shown to “scramble” the spectrum of two-level system populations in superconducting qubits [9].
@@
Here we describe a continuation of the work performed in Ref. [7], using the same four-qubit chip. The charge environment in the qubit substrate responds to ionizing events through multiple physics channels (including e-h pair production, phonon diffusion, and charge trapping) and on time scales ranging from nanoseconds to hours or even days [10, 11, 3]. Accordingly, the qubits in this work are mildly charge-sensitive transmons () that operate as electrometers, each with a sensing area for electric fields in the substrate of hundreds of square microns. The qubit chip was relocated from the Earth’s surface at Madison, WI to the Northwestern EXperimental Underground Site (NEXUS) at Fermilab in Batavia, IL. Previously, correlated jumps in offset charge, associated with gamma ray and cosmic ray impacts, were observed in this qubit array. Underground, over 99 of cosmic ray muons are shielded by the overburden, creating an environment in which the qubit response to gamma radiation can be studied in isolation. We vary the flux of gamma rays incident on the qubit chip and measure the rate and magnitude of ensuing charge burst events via Ramsey tomography. Other energy-resolving detectors operating simultaneously in NEXUS are used to calibrate the flux and spectrum of ionizing radiation. (This study is in some sense complimentary to that of Ref. [8], in which scintillation detectors provide coincidence information between cosmic ray muon events and qubit errors.)
@@
| Shield Open | Shield Closed | Units | |
| Livetime | 23.949 | 22.075 | hours |
| Q1 Rate | mHz | ||
| Q2 Rate | mHz | ||
| Q3 Rate | mHz | ||
| Q4 Rate | mHz | ||
| Average Rate | mHz | ||
| Corrected Rate | mHz | ||
| Calculated Excess Rate | mHz | ||
| Q1-Q2 | Q3-Q4 | Q1-Q3 | Q1-Q4 | Q2-Q3 | Q2-Q4 | Units | |
| Separation | 640 | 340 | 3195 | 3330 | 3180 | 3240 | m |
| Shield Open | mHz | ||||||
| Shield Closed | mHz |
@@
The charge jump rates, averaged across all four qubits, that we measure for the SO and SC data are 0.51 and 0.19 mHz respectively, as shown in Table 1. Therefore, closing the shield reduces the rate of qubit charge bursts by only a factor of 2.7, over seven times less than expected based on the factor of 20 reduction in gamma flux measured in the LMO detector. A possible explanation for this discrepancy is that in our lowest-background SC configuration, we are sensitive to an excess source of charge bursts that is not dominated by the external gamma flux. Potential sources of radiation inside the fridge that might significantly impact the qubit package but not the LMO detector warrant follow up study and assay [17, 18]. The expected muon flux through the qubit (for both configurations) is 0.08 mHz/cm2 – too low to explain the observed rates. The NuMI muon neutrino beam was not active during data collection for this work. Neutrino interactions in matter [19] therefore contributed no additional muon flux during this study.
We infer that the ambient gamma flux does not dominantly contribute to our SC dataset. We therefore subtract the SC rates from the SO data to obtain a reduced burst rate in the SO data, induced by ambient gammas, of mHz. In subtracting these data to determine the gamma-induced component, we account for an unknown population of bursts by assuming a constant excess rate of jumps present in both the SO and SC data. Based on the LMO data (see Supplemental Material), we estimate our ambient gamma flux in the SO configuration to be approximately five times lower than that measured using a NaI detector in Ref. [7]. If we assume, as Ref. [7] does, that the surface burst rate of 1.35 mHz was gamma-dominated, our ambient rate estimate of mHz is proportional with that expected reduction in flux.
The distances between each qubit pair are different, per Table 2 and Figure 2. The smallest separation between qubit pairs (qubits 3 and 4) is 340 m and the largest separation between pairs (qubits 1 and 4) is 3330 m. This variable separation between qubit pairs enables some inferences about correlated noise rates. First, it is technically possible that separate burst events could create conditions that mimic correlated charge jumps arising from a single charge burst. In the SO data, we measure a correlated charge jump rate in nearby qubit pairs of 0.27 mHz for qubits 1 and 2, and 0.29 mHz for qubits 3 and 4. Given the low single-qubit jump rates in Table 1, the rate at which this stochastic coincidence could occur is orders of magnitude lower than the measured correlated jump rate. Next, this correlated rate is approximately half of the single-qubit rate (to within statistical uncertainty), which is consistent with that observed in Ref. [7] when dominated by gamma flux. In the SO data, the correlated jump rate across distant qubit pairs is too low to make any statements on burst origin with any statistical significance. The same statistical limitation is true for nearby qubit pairs in the SC data. Finally, we are able to, for the first time, eliminate correlated charge noise in charge-sensitive qubits separated by over 3 mm, on timescales nearing one day.
@@
The NEXUS DR, shown in Figure S1, is enclosed in a three-part lead shield: (1) a four-inch-thick inner lead plug above the payload, thermalized at 1 K; (2) a four-inch-thick stationary wall; and (3) a cart-mounted, movable lead shield with a nominal thickness of four inches. Taken together, these three shield components provide full, coverage of experimental payloads against ambient gamma radiation. The qubit chip is installed inside of a 1-mm thick A4K hermetic can from Amuneal, thermalized to the DR mixing chamber plate. The DR itself is further equipped with a 1-mm thick A4K magnetic shield from Amuneal, thermalized at 4 K to reduce the magnetic field in the payload region. Finally, a external Metglass blanket surrounds the 300-K shell. The DR base temperature was stabilized at 10.5 mK during the collecation of all data presented here.
@@
Two different warm RF systems are used for qubit characterization and measurement. Both are connected to the qubit input and output ports via a warm RF switch (not shown in Figure S2). A vector network analyzer (not shown in Fig. S2) is used for continuous wave measurements during initial qubit characterization. Pulsed RF measurements are performed using the Quantum Instrumentation and Control Kit (QICK): a Xilinx RFSoC board with custom open-source software and firmware for the control of qubit systems [24, 25]. The qubit charge bias voltage is supplied by an arbitrary waveform generator (AWG). The system has since been upgraded to allow for multiplexed simultaneous readout, but at the time of the measurement presented in the main text, this was not possible. As such, individual qubits are measured consecutively for each bias voltage in a Ramsey tomography scan before moving onto the next voltage value.
@@
The efficiency of this method is assessed by applying it to synthetic data with charge jumps inserted at known locations. For each qubit we simulate 1600 scans with a jump rate of 1.1 mHz (slightly higher than the actual rate to ensure we do not see effects of pile-up). The jumps are randomly injected with sizes selected from a flat probability distribution between 0.01 and 0.5. The simulated scans are produced using the templates discussed above, convolved with a Gaussian noise spectrum to add the characteristic noise of our data. The accuracy of the jump-finding algorithm is gauged by the fraction of known charge jumps in the synthetic data that are correctly identified. Systematic errors on this quantity are derived from the standard deviation of this value for 15 different jump sizes across 75 sets of simulated scans per qubit (see Table S2). This effectively captures the variation in efficiency across the range of , which can be crudely approximated by Gaussian (with slightly lower efficiency near the threshold in particular).
@@
We exploit the measured ratio between the SO and SC LMO datasets by solving for a constant excess charge jump rate present in both datasets and the gamma-induced charge jump rate for each. By solving,
where is the measured charge jump rate in the SO and SC data and is corrected gamma-induced rate in each dataset, we find an excess charge jump rate of . These results are presented in Table 1 of the main text.